OpenStudy (tmcluv):

Which of the following ordered pairs represents the solution to the system given below? 2x + y = 20 4x − 2y = 40 (10, 0) (0, 10) (4, 2) (10, −10)

OpenStudy (zyi6):

Let's plug in the first answer choice to see if it works: 2(10) + (0) = 20 20 + 0 = 20 20 = 20 Works here! 4(10) - 2(0) = 40 40 - 0 = 40 40 = 40 Works here too! So, (10, 0) is your answer!